The equivalent weight of an divalent metal whose chloride has vapour density equal to 60 is
1)30 2)49 3)24.5 4)16.5

Grade:11

1 Answers

Rayon
5 years ago
Cool and nice question. Liked it. 
 
From given information as it is divalent metal it combines with chlorine like MCl².
 
As per given vapour density is 60. So molecular weight is 120. Now as we know molecular weight of chlorine and that of MCl².
 
Finding atomic weight of M. So atomic weight of M is 120 - 2(35.5) = 49.
 
Equivalent weight  =  (Molecular weight) /(Valence)  =  49/2 = 24.5.